# Zero-Shot Skeleton-based Action Recognition with Dual Visual-Text Alignment

### v1

@@ Abstract
Zero-shot action recognition, which addresses the issue of scalability and generalization in action recognition and allows the models to adapt to new and unseen actions dynamically, is an important research topic in computer vision communities. The key to zero-shot action recognition lies in aligning visual features with semantic vectors representing action categories. Most existing methods either directly project visual features onto the semantic space of text category or learn a shared embedding space between the two modalities. However, a direct projection cannot accurately align the two modalities, and learning robust and discriminative embedding space between visual and text representations is often difficult. To address these issues, we introduce Dual Visual-Text Alignment (DVTA) for skeleton-based zero-shot action recognition. The DVTA consists of two alignment modules—Direct Alignment (DA) and Augmented Alignment (AA)—along with a designed Semantic Description Enhancement (SDE). The DA module maps the skeleton features to the semantic space through a specially designed visual projector, followed by the SDE, which is based on cross-attention to enhance the connection between skeleton and text, thereby reducing the gap between modalities. The AA module further strengthens the learning of the embedding space by utilizing deep metric learning to learn the similarity between skeleton and text. Our approach achieves state-of-the-art performances on several popular zero-shot skeleton-based action recognition benchmarks.
Index Terms:
Zero-Shot Action Recognition, Skeleton-Based Action Recognition.
I Introduction
@@
Action recognition emerges as a hot topic in research due to its wide applications in areas such as video understanding [1, 2], human-computer interaction [3], health monitoring [4] and robotics. However, with the rapid increase in the number of videos generated due to the prevalence of social media, smartphones, and surveillance cameras, collecting video action labels becomes increasingly impractical and burdensome. Consequently, Zero-Shot Action Recognition (ZSAR) [5] emerges as a viable solution. ZSAR alleviates the burden of collecting labeled data by leveraging a shared semantic space to generalize knowledge across unseen action categories, thus enabling recognition of actions without requiring explicit training examples.
@@
The key to the implementation of ZSAR lies in effectively utilizing the shared semantic space to achieve recognition of unseen classes. A common approach is the projection-based method. For instance, the early method [6] projects low-level visual features into a semantic space composed of word embeddings. However, these methods solely relying on direct visual projection struggle to bridge the gap between modalities. Another typical way is to generate an intermediate space for visual-semantic fusion. For example, BiDiLEL [7] employs two-stage mapping, separately mapping visual and semantic features into a latent space. These approaches face the challenge of designing an effective embedding space that can capture the nuances and complexities of visual and text modalities. Moreover, the above methods utilize RGB video data, which is susceptible to background interference, leading to suboptimal performance in ZSAR.
@@
With the emergence of depth sensors and advancements in pose estimation algorithms [8], skeleton-based action recognition has gained attention. Skeleton data is robust to lighting and environmental changes, making it a good alternative to RGB video data. However, compared to RGB video data, there is relatively less research on ZSAR using skeleton data. Recent methods perform distribution alignment to enhance modality correspondence. To capture complex correlations between the distributions of the visual space and the text semantic space, SMIE [9] addresses skeleton-based ZSAR via mutual information estimation and maximization. However, these methods often overlook the significant differences between the two spaces and directly use connection models on the raw extracted features to capture dependencies, which limits their performance to a certain extent.
@@
Specifically, we propose the Dual Visual-Text Alignment (DVTA) network, which employs two key components for modality alignment. The first component is Direct Alignment (DA), which utilizes a specially designed deep neural network for visual projection to map skeleton features into the semantic space. An attention-based Semantic Description Enhancement (SDE) module is then used to integrate visual information to text features, which establishes a closer connection between skeleton sequences and text descriptions. During training, a supervised training loss is formulated to generate more positive examples, maximizing the similarity between the paired skeleton and text features within this space. The second component is the Augmented Alignment (AA) module, which uses a neural network to further align distributions by learning similarity scores between different modality features within the latent space. The LeakySigmoid activation function is designed to normalize the predicted similarity values. The KL divergence loss is employed to jointly optimize the alignment modules, enhancing the alignment between the two modalities.
@@
  • We propose an Dual Visual-Text Alignment (DVTA), a novel zero-shot approach for skeleton-based action recognition. The method enhances generalization to unseen classes by jointly optimizing two modules: Direct Alignment (DA) and Augmented Alignment (AA).
@@
Figure 2: Framework of the proposed Dual Visual-Text Alignment (DVTA). It jointly optimizes two alignment modules—Direct Alignment (DA) and Augmented Alignment (AA)—in a supervised manner. The DA module maps skeleton features and text features to a common latent space using different projection modules. Subsequently, the Semantic Description Enhancement (SDE) submodule, which utilizes the cross-attention mechanism, generates augmented text features containing visual information, thereby fostering a closer connection between the skeleton sequence and the text description. The AA module computes the correlation between the two modalities using a designed Deep Metric Network (DMN) and further aligns their distributions by normalizing the similarity scores through the LeakySigmoid activation function. The KL divergence loss incorporating a greater number of positive examples is employed as the contrastive loss to jointly optimize the two alignment modules, thereby enhancing the alignment between the two modalities.
II Related Work
II-A Zero-Shot Action Recognition
@@
With the rapid development of large-scale language models, new categories can be recognized using a shared semantic space, leading to an increasing interest in zero-shot action recognition. Currently, most approaches are dedicated to aligning the visual space [10, 7] and the semantic space [11, 12]. Numerous studies [13, 14, 15, 16, 11] construct an end-to-end framework to achieve the transformation from the original visual input to the final visual semantic representation in a unified way. In order to capture inter-modal correlations more accurately, attentional mechanisms are designed to focus on learning the multimodal space [17, 18]. In addition, augmented category information methods [19, 20] can further improve the performance of zero-shot learning. Recently, DeCalGAN [21] generates unseen action video features to transfer knowledge from known action categories to the feature distribution of unknown categories.
II-B Skeleton-Based Action Recognition
@@
We primarily review methods based on Graph Convolutional Network (GCN), which are used to extract skeleton features. Yan et al. [22] propose the innovative ST-GCN, which represents human joints as graph structures and applies spatio-temporal GCN for the first time to skillfully capture the spatial and temporal features of the human skeleton. This pioneering work leads to the thriving development of GCN-based skeleton action recognition methods [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Li et al. [23] introduce the actional-structural graph convolution network (AS-GCN), which generates a generalized skeleton graph to capture richer dependencies between joints. Shi et al. [24] propose 2s-AGCN, which utilizes second-order information of skeleton data to enhance recognition accuracy and dynamically learns graph topology to improve the model’s generality. To further improve efficiency, Cheng et al. [25, 26] propose the shift graph convolutional network (Shift-GCN) to significantly reduce the computational complexity of the GCN method. Subsequently, they also introduce a more lightweight Shift-GCN++ [26] to further optimize performance. Chen et al. [27] develop a Channel-wise Topology Refinement Graph Convolution (CTR-GCN) to dynamically learn the topology and capture the inter-node relationships in a more fine-grained manner. Chi et al. [28] adopt an information bottleneck to guide the model in learning latent representations, and introduce attention-based graph convolution to capture contextual intrinsic topology. Recently, Xiang et al. [29] propose a Generative Action-description Prompts (GAP) approach, which utilizes action semantic information to enhance the representation learning ability of the skeleton encoder. These research works achieve satisfactory performance in skeleton feature extraction, providing strong support for encoders in the field of skeleton-based action recognition.
II-C Zero-Shot Skeleton-Based Action Recognition
@@
Aligning the skeleton space with the semantic space without accessing unseen class samples for training poses a significant challenge, resulting in relatively few works on zero-shot skeleton action recognition. Zero-shot learning methods previously applied to other domains are creatively extended to this domain. For example, DeViSE [34], originally designed for visual image classification tasks, maps extracted skeleton features to the text feature space through a learnable linear projection. ReViSE [35] employs the maximum average discrepancy loss function to further facilitate the alignment between the two modalities. RelationNet [36] calculates the similarity between the projected features using the attribute network and the relation network. Subsequently, JPoSE [37] proposes the construction of a separate multimodal space for each PoS tag to complete cross-modal fine-grained action retrieval. CADA-VAE [38] utilizes a variational autoencoder to learn the latent space between modalities. More recently, works specifically targeting zero-shot skeleton-based action recognition have emerged. For instance, SynSE [39] leverages syntactic structure partition labels to align skeleton features with corresponding verb embeddings and noun embeddings. SMIE [9] aligns the distributions of the two modalities by maximizing mutual information. Some works utilize large-scale language models to generate extensive semantic information for modality alignment. For instance, PURLS [40] and STAR [41] employ large-scale language models to create multi-part descriptions for fine-grained alignment, while InfoCPL [42] combines numerous descriptive texts to enrich semantic information, improving recognition performance. However, these approaches rely heavily on manual preparation of extensive text information, which is impractical for real-world applications. Our approach differs from these methods in two key ways. First, we focus on structural innovation by designing a specialized deep visual projector to bridge the substantial gap between skeleton and text features. Additionally, we propose a dual alignment strategy that jointly optimizes two alignment networks to strengthen the correlation between skeleton and text features. Second, we enrich semantic information using only a minimal set of label descriptions, significantly reducing the need for manual preparation.
III Method
@@
Zero-shot action recognition aims to recognize unseen action categories in the test set, despite the absence of corresponding training skeleton samples. The key of achieving this lies in aligning the visual feature space with the text semantic space, leveraging the joint embedding space to recognize the action of samples from unknown categories. Nevertheless, aligning visual and text remains a challenging task due to the semantic gap between the skeleton sequence and the corresponding text category. To bridge this gap, we introduce a Dual Visual-Text Alignment (DVTA), which mainly comprises two alignment modules, Direct Alignment (DA) and Augmented Alignment (AA). Figure 2 illustrates the overall structure of the proposed method.
III-A Visual-Text Embedding
@@
III-A1 Deep Feature Projection
@@
The features $v$ and $t$ come from different spaces and the dimensions are also different. Two projection networks are designed to map the features of skeleton and text to a common embedding space. In order to map the skeleton features to the same dimension as the text features to a greater extent, a depth nonlinear mapping projector is used to transform the skeleton features to the common embedding space, the effect of which will be reflected in section IV-D. For features of text, a linear layer is utilized to map them to the common space. Let $F_{v}$ and $F_{t}$ be the projectors for skeleton features and text features, respectively. To simplify the expressions, the superscripts for visible and unseen classes are omitted. The transformed skeleton and text features are:
@@
III-A2 Semantic Description Enhancement
@@
In zero-shot action recognition, the accuracy and richness of semantic space information are critical for the generalization to unseen samples. The inspiration for semantic description enhancement comes from the complex process in the human brain, where both action relevance and environmental context are considered during action recognition. By enriching the semantic content of actions through context augmentation, the model’s recognition capability can be improved. We use GPT-3 [43] to generate contextual descriptions for action labels by employing a fixed prompt template: “Describe the context in which the $<$action$>$ occurs”. For example, for the action “eat meal”, a descriptive sentence might be: “The person is sitting at a table, facing a plate of food in front of them.”. This context description, after processing through the above text feature extraction operations, yields the action context text feature $t_{cont}$. We base on a cross-attention mechanism to jointly embed the visual feature $v_{e}$, label text feature $t_{e}$, and context text feature $t_{cont}$ to generate the semantically augmented text feature $t_{aug}$. Specifically, we define $K$ and $V$ as $T=stack({t_{e}},{t_{cont}})\in{\mathbb{R}^{2\times h}}$, and $Q=v$ as the visual query. The semantically augmented text feature $t_{aug}$ can be obtained as:
@@
III-B Dual Alignment Network
@@
III-B1 Direct Alignment
Given that the number of skeleton feature samples significantly exceeds the number of action labels, it is reasonable to expect that a batch of data will contain a large number of skeleton samples corresponding to a single label, indicating the presence of multiple positive examples. Therefore, the cross-entropy loss is inappropriate. Instead, we utilize KL divergence as the skeleton-text contrastive loss, allowing each skeleton feature sample to form multiple positive examples with the corresponding same-label text features, as referenced in [44, 29]. The similarity score between the skeleton features and text features can be calculated as:
@@ III-B1 Direct Alignment
III-B2 Augmented Alignment
We first concatenate the projected skeleton feature $v$ and text feature $t$, then feed the concatenated feature to the Deep Metric Network $E$ to predict similarity score $G({v},{t})$ between the two features:
$G({v},{t})=LeakySigmoid(E({v},{t})),$ (6)

### v2

@@ Abstract
Zero-shot action recognition, which addresses the issue of scalability and generalization in action recognition and allows the models to adapt to new and unseen actions dynamically, is an important research topic in computer vision communities. The key to zero-shot action recognition lies in aligning visual features with semantic vectors representing action categories. Most existing methods either directly project visual features onto the semantic space of text category or learn a shared embedding space between the two modalities. However, a direct projection cannot accurately align the two modalities, and learning robust and discriminative embedding space between visual and text representations is often difficult. To address these issues, we introduce Dual Visual-Text Alignment (DVTA) for skeleton-based zero-shot action recognition. The DVTA consists of two alignment modules—Direct Alignment (DA) and Augmented Alignment (AA)—along with a designed Semantic Description Enhancement (SDE). The DA module maps the skeleton features to the semantic space through a specially designed visual projector, followed by the SDE, which is based on cross-attention to enhance the connection between skeleton and text, thereby reducing the gap between modalities. The AA module further strengthens the learning of the embedding space by utilizing deep metric learning to learn the similarity between skeleton and text. Our approach achieves state-of-the-art performances on several popular zero-shot skeleton-based action recognition benchmarks. The code is available at: https://github.com/jidongkuang/DVTA.
keywords:
Zero-Shot Action Recognition, Skeleton-Based Action Recognition
journal: Pattern Recognition
1 Introduction
@@
Action recognition emerges as a hot topic in research due to its wide applications in areas such as video understanding, human-computer interaction, health monitoring and robotics. However, with the rapid increase in the number of videos generated due to the prevalence of social media, smartphones, and surveillance cameras, collecting video action labels becomes increasingly impractical and burdensome. Consequently, Zero-Shot Action Recognition (ZSAR) [1] emerges as a viable solution. ZSAR alleviates the burden of collecting labeled data by leveraging a shared semantic space to generalize knowledge across unseen action categories, thus enabling recognition of actions without requiring explicit training examples.
@@
The key to implementation of ZSAR lies in effectively utilizing the shared semantic space to achieve recognition of unseen classes. A common approach is the projection-based method. For instance, the early method [2] projects low-level visual features into a semantic space composed of word embeddings. However, these methods solely relying on direct visual projection struggle to bridge the gap between modalities. Another typical way is to generate an intermediate space for visual-semantic fusion. For example, BiDiLEL [3] employs two-stage mapping, separately mapping visual and semantic features into a latent space. These approaches face the challenge of designing an effective embedding space that can capture the nuances and complexities of visual and text modalities. Moreover, the above methods utilize RGB video data, which is susceptible to background interference, leading to suboptimal performance in ZSAR.
@@
With the emergence of depth sensors and advancements in pose estimation algorithms, skeleton-based action recognition [4, 5] has gained attention. Skeleton data is robust to lighting and environmental changes, making it a good alternative to RGB video data. However, compared to RGB video data, there is relatively less research on ZSAR using skeleton data. Recent methods perform distribution alignment to enhance modality correspondence. To capture complex correlations between the distributions of the visual space and the text semantic space, SMIE [6] addresses skeleton-based ZSAR via mutual information estimation and maximization. However, these methods often overlook the significant differences between the two spaces and directly use the connection models on the raw extracted features to capture dependencies, which limits their performance to a certain extent.
@@
Specifically, we propose the Dual Visual-Text Alignment (DVTA) network, which employs two key components for modality alignment. The first component is Direct Alignment (DA), which utilizes a specially designed deep neural network for visual projection to map skeleton features into the semantic space. An attention-based Semantic Description Enhancement (SDE) module is then used to integrate visual information into text features, which establishes a closer connection between skeleton sequences and text descriptions. During training, a supervised training loss is formulated to generate more positive examples, maximizing the similarity between the paired skeleton and text features within this space. The second component is the Augmented Alignment (AA) module, which uses a neural network to further align distributions by learning similarity scores between different modality features within the latent space. The LeakySigmoid activation function is designed to normalize the predicted similarity values. The KL divergence loss is employed to jointly optimize the alignment modules, enhancing the alignment between the two modalities.
@@
  • 1.
    We propose a Dual Visual-Text Alignment (DVTA), a novel zero-shot approach for skeleton-based action recognition. The method enhances generalization to unseen classes by jointly optimizing two modules: Direct Alignment (DA) and Augmented Alignment (AA).
  • 2.
@@
2 Related Work
2.1 Zero-Shot Action Recognition
@@
With the rapid development of large-scale language models, new categories can be recognized using a shared semantic space, leading to an increasing interest in zero-shot action recognition. Currently, most approaches are dedicated to aligning the visual space [7] and the semantic space [8]. Numerous studies [9, 8] construct an end-to-end framework to achieve the transformation from the original visual input to the final visual semantic representation in a unified way. In order to capture inter-modal correlations more accurately, attentional mechanisms are designed to focus on learning the multimodal space [10]. In addition, augmented category information methods [11] can further improve the performance of zero-shot learning. Recently, DeCalGAN [12] generates unseen action video features to transfer knowledge from known action categories to the feature distribution of unknown categories.
2.2 Skeleton-Based Action Recognition
@@
We primarily review methods based on Graph Convolutional Network (GCN), which are used to extract skeleton features. Yan et al. [13] propose the innovative ST-GCN, which represents human joints as graph structures and applies spatio-temporal GCN for the first time to skillfully capture the spatial and temporal features of the human skeleton. This pioneering work leads to the thriving development of GCN-based skeleton action recognition methods [14, 15, 16]. Li et al. [17] introduce the actional-structural graph convolution network (AS-GCN), which generates a generalized skeleton graph to capture richer dependencies between joints. Shi et al. [18] propose 2s-AGCN, which utilizes second-order information of skeleton data to enhance recognition accuracy and dynamically learns graph topology to improve the model’s generality. To further improve efficiency, Cheng et al. [19, 20] propose the shift graph convolutional network (Shift-GCN) to significantly reduce the computational complexity of the GCN method. Subsequently, they also introduce a more lightweight Shift-GCN++ [20] to further optimize performance. Chen et al. [21] develop a Channel-wise Topology Refinement Graph Convolution (CTR-GCN) to dynamically learn the topology and capture the inter-node relationships in a more fine-grained manner. Chi et al. [22] adopt an information bottleneck to guide the model in learning latent representations, and introduce attention-based graph convolution to capture contextual intrinsic topology. Recently, Xiang et al. [23] propose a Generative Action-description Prompts (GAP) approach, which utilizes action semantic information to enhance the representation learning ability of the skeleton encoder. [24] and [25] use the multi-granularity information of skeleton to improve recognition ability. These research works achieve satisfactory performance in skeleton feature extraction, providing strong support for encoders in the field of skeleton-based action recognition.
2.3 Zero-Shot Skeleton-Based Action Recognition
@@
Aligning the skeleton space with the semantic space without accessing unseen class samples for training poses a significant challenge, resulting in relatively few works on zero-shot skeleton action recognition. Zero-shot learning methods previously applied to other domains are creatively extended to this domain. For example, DeViSE [26], originally designed for visual image classification tasks, maps extracted skeleton features to the text feature space through a learnable linear projection. ReViSE [27] employs the maximum average discrepancy loss function to further facilitate the alignment between the two modalities. RelationNet [28] calculates the similarity between the projected features using the attribute network and the relation network. Subsequently, JPoSE [29] proposes the construction of a separate multimodal space for each PoS tag to complete cross-modal fine-grained action retrieval. CADA-VAE [30] utilizes a variational autoencoder to learn the latent space between modalities. More recently, works specifically targeting zero-shot skeleton-based action recognition have emerged. For instance, SynSE [31] leverages syntactic structure partition labels to align skeleton features with corresponding verb embeddings and noun embeddings. SMIE [6] aligns the distributions of the two modalities by maximizing mutual information. Some works utilize large-scale language models to generate extensive semantic information for modality alignment. For instance, PURLS [32] and STAR [33] employ large-scale language models to create multi-part descriptions for fine-grained alignment, while InfoCPL [34] combines numerous descriptive texts to enrich semantic information, improving recognition performance. However, these approaches rely heavily on the manual preparation of extensive text information, which is impractical for real-world applications. Our approach differs from these methods in two key ways. First, we focus on structural innovation by designing a specialized deep visual projector to bridge the substantial gap between skeleton and text features. Additionally, we propose a dual alignment strategy that jointly optimizes two alignment networks to strengthen the correlation between skeleton and text features. Second, we enrich semantic information using only a minimal set of label descriptions, significantly reducing the need for manual preparation.
3 Method
@@
Zero-shot action recognition aims to recognize unseen action categories in the test set, despite the absence of corresponding training skeleton samples. The key to achieving this lies in aligning the visual feature space with the text semantic space, leveraging the joint embedding space to recognize the action of samples from unknown categories. Nevertheless, aligning visual and text remains a challenging task due to the semantic gap between the skeleton sequence and the corresponding text category. To bridge this gap, we introduce a Dual Visual-Text Alignment (DVTA), which mainly comprises two alignment modules, Direct Alignment (DA) and Augmented Alignment (AA).
Figure 2 illustrates the overall structure of the proposed method. Specifically, the DA module maps skeleton features and text features to a common latent space using different projection modules. Subsequently, the Semantic Description Enhancement (SDE) submodule, which utilizes the cross-attention mechanism, generates augmented text features containing visual information, thereby fostering a closer connection between the skeleton sequence and the text description. The AA module computes the correlation between the two modalities using a designed Deep Metric Network (DMN) and further aligns their distributions by normalizing the similarity scores through the LeakySigmoid activation function. The KL divergence loss incorporating a greater number of positive examples is employed as the contrastive loss to jointly optimize the two alignment modules, thereby enhancing the alignment between the two modalities.
3.1 Visual-Text Embedding
@@
Figure 2: Framework of the proposed Dual Visual-Text Alignment (DVTA). It jointly optimizes two alignment modules—Direct Alignment (DA) and Augmented Alignment (AA)—in a supervised manner.
3.1.1 Deep Feature Projection
@@
The features $v$ and $t$ come from different spaces and the dimensions are also different. Two projection networks are designed to map the features of skeleton and text to a common embedding space. In order to map the skeleton features to the same dimension as the text features to a greater extent, a learnable depth nonlinear mapping projector is used to transform the skeleton features to the common embedding space, the effect of which will be reflected in section 4.4. A linear layer is utilized to map text features to the common space. Let $F_{v}$ and $F_{t}$ be the projectors for skeleton features and text features, respectively. To simplify the expressions, the superscripts for visible and unseen classes are omitted. The transformed skeleton and text features are:
@@
3.1.2 Semantic Description Enhancement
@@
In zero-shot action recognition, the accuracy and richness of semantic space information are critical for the generalization to unseen samples. The inspiration for semantic description enhancement comes from the complex process in the human brain, where both action relevance and environmental context are considered during action recognition. By enriching the semantic content of actions through context augmentation, the model’s recognition capability can be improved. We use GPT-3 [35] to generate contextual descriptions for action labels by employing a fixed prompt template: “Describe the context in which the <<action>> occurs”. For example, for the action “eat meal”, a descriptive sentence might be: “The person is sitting at a table, facing a plate of food in front of them.”. This context description, after processing through the above text feature extraction operations, yields the action context text feature $t_{cont}$. We base on a cross-attention mechanism to jointly embed the visual feature $v_{e}$, label text feature $t_{e}$, and context text feature $t_{cont}$ to generate the semantically augmented text feature $t_{aug}$. Specifically, we define $K$ and $V$ as $T=stack({t_{e}},{t_{cont}})\in{\mathbb{R}^{2\times h}}$, and $Q=v$ as the visual query. The semantically augmented text feature $t_{aug}$ can be obtained as:
@@
3.2 Dual Alignment Network
@@
3.2.1 Direct Alignment
Given that the number of skeleton feature samples significantly exceeds the number of action labels, it is reasonable to expect that a batch of data will contain a large number of skeleton samples corresponding to a single label, indicating the presence of multiple positive examples. Therefore, the cross-entropy loss is inappropriate. Instead, we utilize KL divergence as the skeleton-text contrastive loss, allowing each skeleton feature sample to form multiple positive examples with the corresponding same-label text features, as referenced in [36, 23]. The similarity score between the skeleton features and text features can be calculated as follows:
@@ 3.2.1 Direct Alignment
3.2.2 Augmented Alignment
We first concatenate the projected skeleton feature $v$ and text feature $t$, then feed the concatenated feature to the Deep Metric Network $E$ to predict similarity score G(v,t)G({v},{t}) between the two features:
G(v,t)=LeakySigmoid(E(v,t)),G({v},{t})=LeakySigmoid(E({v},{t})), (6)